OpenStudy (anonymous):

Simplify and select the answer with the appropriate restrictions for the variables. 6x^2-3x -------- 3x^2

OpenStudy (anonymous):

a. 2x-1, (x does not = 0) b. 2x-1/x, (x does not =0) c. 2-3x, (x does not = 0) d. 6x-1, (x does not = 0)